OpenStudy (anonymous):

Four times the sum of a number and 15 is at least 120. Let x represent the number. Find all possible values for x. A. x ≥ 26 B. x ≥ –15 C. x ≥ 15 D. x ≥ –26

OpenStudy (gorv):

let number is x 4*x+15>=120 4x+15>=120 now solve it ..give it a try

OpenStudy (imstuck):

@gorv you are not correct in that inequality statement.

OpenStudy (imstuck):

"Four times the sum of a number and 15" is actually written like this: 4(x+15). And the sign is a greater than or equal to sign, not an equals sign.

OpenStudy (anonymous):

C?

OpenStudy (imstuck):

\[4(x+15)\ge 120\]Now try to solve for x.

OpenStudy (imstuck):

yes, good job!

OpenStudy (gorv):

i skipped a

OpenStudy (gorv):

thanks by the way for correcting

jhonyy9 (jhonyy9):

4(x+15)>=120 4x+60>=120 can you continue it ?

jhonyy9 (jhonyy9):

so 4x>=60 divide both sides by 4 x>= 15 ok ?